OpenStudy (anonymous):

A company sells boxes of cow bells (b) for $20 and boxes of air horns (h) for $30. They can make a batch of cow bells that fills 8 boxes or a batch of air horns that fills 9 boxes. The company only has 53 boxes. Which system of equations helps the company plan to make $275?

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

5x - 3(15x+6) = -13 -40x=-13+18 x=-1/8 y=-15/8+6=33/8

OpenStudy (anonymous):

A. b = 53 - h 8b + 9h = 53 B. 8b = 53 - 9h 20b + 30h = 275 C. b = 53 - h 20b + 30h = 275 D. 8b = 53 - 9h 20b - 30h = 275 Which one does that correspond to? And how'd you get those numbers? D:

OpenStudy (texaschic101):

8b + 9h = 53 --> 8b = 53 - 9h 20b + 30h = 275

OpenStudy (anonymous):

Okay, so 8b is bells, 9h is horns? And then the second equation was 20 for the boxes of bells and 30 for the boxes of horns. Right?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

Okay thanks! I have a similar one, but it doesn't make since to me. Captain Stacey is going to load her sailboat with jars of honey. She has room for 11 jars. She has allotted for 58 pounds of honey on her boat. She can choose from 3-pound jars (the variable j) or 7-pound jars (the variable h). Which equations would be a constraint to Captain Stacey maximizing her honey load? a. j - h = 11 b.j = 11 + h c.3j = 58 + 7h d.3j = 58 - 7h

OpenStudy (anonymous):

(I took this homework/quiz thing before and these are the ones I don't understand.)

OpenStudy (ranga):

If she chooses "j" number of 3-pound jars and "h" number of 7-pound jars the total weight will be how much?

OpenStudy (texaschic101):

I am thinking D. But I have a little trouble with the constraints, so you need to get a second opinion

OpenStudy (ranga):

yes d) is what I am getting too.

OpenStudy (texaschic101):

you continue with your explaining ranga....I do not know how to explain this because I have problems with it

OpenStudy (anonymous):

11, because she only has room for eleven jars?

OpenStudy (ranga):

If she loads "j" number of 3-pound jars and "h" number of 7-pound jars on the sailboat the total weight will be: 3j + 7h pounds. She has allotted 58 pounds for honey and so 3j + 7h = 58 is one constraint The second constraint is: j + h = 11 and this constraint is not covered in the answer choices. So we have to go with the first constraint 3j + 7h = 58

OpenStudy (anonymous):

Oh! I get it. Thanks!

OpenStudy (texaschic101):

that helped me too....thanks ranga :)

OpenStudy (ranga):

you are all welcome.

OpenStudy (texaschic101):

so a constraint is just a limitation ?

OpenStudy (texaschic101):

nvm...I understand :)

OpenStudy (ranga):

Here the second constraint will be better stated as: j + h <= 11 (instead of just the equal sign). Just because she has space for 11 jars does not mean all space must be used. But the first constraint of 3j + 7h = 58 should be an equal sign because she wants to maximize the mount of honey she carries on the sailboat.